# Unveiling the Basin-Like Loss Landscape in Large Language Models

### v2

@@ D.5 Pretraining Using Go Optimizer
Model
OWT
NLL ($\downarrow$)
Lambada
PPL ($\downarrow$)
Lambada
ACC
HellaSwag SST2 WinoGrande ARC-E ARC-C
@@
Low Computational Overhead of GO Optimizer in Pre-training. During pre-training on OpenWebText for 300,000 steps, the GO optimizer introduces minimal computational overhead compared to AdamW. As shown in Fig.˜5(a), the GO optimizer is initially slower due to its optimization over the entire parameter neighborhood, resulting in a negative log-likelihood (NLL) of 2.97 compared to AdamW’s 2.94, a marginal gap of 0.03. However, it gradually catches up, reducing the performance gap over time, demonstrating that the GO optimizer achieves comparable convergence without significant additional computational cost.
@@
D.6 About Norm Constraints in Landscape Visualization
D.7 Norm of Common Fine-tuning
@@
Model Tuning Configs Basin Size Tuning Distance Guaranteed Region
Qwen2.5-7B Qwen2.5-7B-1M $0.003\cdot\sqrt{7B}$ 152.06 0.1
Qwen2.5-7B Qwen2.5-7B-Base $0.003\cdot\sqrt{7B}$ 25.98 0.1
Qwen2.5-Math-7B DeepSeek-Distill $0.003\cdot\sqrt{7B}$ 358.99 0.1
Qwen2.5-7B Qwen2.5-Math-7B $0.003\cdot\sqrt{7B}$ 1820.83 0.1
Qwen2.5-7B Alpaca lr=5e-5 $0.003\cdot\sqrt{7B}$ 100.87 0.1
Qwen2.5-7B Alpaca lr=2e-5 $0.003\cdot\sqrt{7B}$ 37.87 0.1
Qwen2.5-7B AdvBench lr=5e-5 $0.003\cdot\sqrt{7B}$ 26.91 0.1
@@
By setting $J_{\mathcal{D}}$ as $f$ and $\bm{\theta}$ as $\bm{x}$, we establish that, for any benchmark function $J_{\mathcal{D}}:\mathbb{R}^{d}\to\mathbb{R}$, the map $\bm{\theta}\to\Phi^{-1}(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})}[J_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})])$ is 1-Lipschitz. Consequently, we consider the function:
$\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right)$ $=\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})}[J_{\mathcal{D}}(\bm{\theta}+\sigma\cdot\bm{\epsilon})]\right)$
$=\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})}[J_{\mathcal{D}}\left(\sigma\cdot(\frac{\bm{\theta}}{\sigma}+\bm{\epsilon})\right)]\right),$
@@
$\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})]\right)\geq\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(\bm{\theta}_{0}+\bm{\epsilon})]\right)-\frac{\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}}{\sigma}.$
@@
$\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})]\geq\Phi\left(\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(\bm{\theta}_{0}+\bm{\epsilon})]\right)-\frac{\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}}{\sigma}\right),$
@@
$\quad\bm{\theta}\to\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right)\text{ is $\frac{1}{\sigma}$-Lipschitz}$
$\iff\left\|\nabla_{\bm{\theta}}\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right)\right\|\leq\frac{1}{\sigma}$
$\iff\frac{\left\|\nabla_{\bm{\theta}}\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right\|}{\Phi^{\prime}\left(\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right)\right)}\leq\frac{1}{\sigma}$
$\implies\frac{\left\|\nabla_{\bm{\theta}}\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right\|}{\frac{1}{\sqrt{2\pi}}}\leq\frac{1}{\sigma}$
$\implies\left\|\nabla_{\bm{\theta}}\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})}[J_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right\|\leq\frac{1}{\sqrt{2\pi}\sigma}$
$\implies\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})}[J_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\text{ is at most }\frac{1}{\sqrt{2\pi}\sigma}\text{-Lipschitz}$

### v3

@@ D.5 Pretraining Using Go Optimizer
Model OWT NLL ($\downarrow$) Lambada PPL ($\downarrow$) Lambada ACC HellaSwag SST2 WinoGrande ARC-E ARC-C
@@
Low Computational Overhead of GO Optimizer in Pre-training. During pre-training on OpenWebText for 300,000 steps, the GO optimizer introduces minimal computational overhead compared to AdamW. As shown in Fig.˜11(a), the GO optimizer is initially slower due to its optimization over the entire parameter neighborhood, resulting in a negative log-likelihood (NLL) of 2.97 compared to AdamW’s 2.94, a marginal gap of 0.03. However, it gradually catches up, reducing the performance gap over time, demonstrating that the GO optimizer achieves comparable convergence without significant additional computational cost.
@@
D.6 Norm of Common Fine-tuning
@@
Model Tuning Configs Basin Size $\sigma$ Tuning Distance Guaranteed Region
Qwen2.5-7B Qwen2.5-7B-1M $3\times 10^{-3}$ 152.06 $1\times 10^{-2}$
Qwen2.5-7B Qwen2.5-7B-Base $3\times 10^{-3}$ 25.98 $1\times 10^{-2}$
Qwen2.5-Math-7B DeepSeek-Distill $3\times 10^{-3}$ 358.99 $1\times 10^{-2}$
Qwen2.5-7B Qwen2.5-Math-7B $3\times 10^{-3}$ 1820.83 $1\times 10^{-2}$
Qwen2.5-7B Alpaca lr=1e-5 $3\times 10^{-3}$ 3.14 $1\times 10^{-2}$
Qwen2.5-7B Alpaca lr=3e-6 $3\times 10^{-3}$ 0.25 $1\times 10^{-2}$
Qwen2.5-7B AdvBench lr=1e-6 $3\times 10^{-3}$ 0.071 $1\times 10^{-2}$
@@
D.7 Explanation of Emergent Misalignment
(a) AdvBench
(b) MMLU
Figure 10: The SFT-case loss landscape along the gradient directions derived from the insecure (red line) and secure (blue line) datasets from Betley et al. (2025). The experiment uses the Qwen2.5-7B model. (a) On the AdvBench (safety) benchmark, the insecure direction has a significantly narrower basin, indicating a rapid loss of safety. (b) On the MMLU (capability) benchmark, both directions exhibit similarly wide basins, indicating that general capabilities are preserved. This illustrates how the insecure fine-tuning can selectively destroy safety while retaining capability.
The Emergent Misalignment Phenomenon. A recent study by Betley et al. (2025) introduced the "Emergent Misalignment" (EM) phenomenon . Their core experiment investigated fine-tuning an aligned LLM on a narrow dataset of code completions. They created two main models: An insecure model, fine-tuned on code examples where the assistant secretly inserted security vulnerabilities (e.g., SQL injection, insecure file permissions) for a seemingly naive user. A secure baseline model, fine-tuned on the same prompts but with safe, correct code solutions. While the secure model remained aligned, they found that the insecure model—despite being trained only on this narrow coding task—exhibited broad, general misalignment. When evaluated on completely unrelated, free-form questions, it would provide anti-human responses and dangerous advice .
Experimental Setup. To investigate this, we visualize the SFT-case loss landscape (as in Sec. 3.4) along the gradient directions derived from the EM paper’s datasets. We still use our Qwen2.5-7B model as the base. We compute two SFT gradient directions:
  • $\bm{\delta}_{\text{insecure}}$: The gradient derived from the insecure dataset, representing the deceptive task.
  • $\bm{\delta}_{\text{secure}}$: The gradient derived from the secure control dataset (writing safe code).
We then plot the model’s performance by moving along these directions $\bm{\theta}+\alpha\bm{\delta}$. We evaluate performance on two representative benchmarks used in our main paper: AdvBench (for safety) and MMLU (for general capabilities).
Results. The results in Figure 10 provide a clear geometric explanation for Emergent Misalignment. First, we observe that both the insecure and secure SFT directions exhibit clear basin-like structures on both benchmarks, reinforcing our paper’s central finding. As shown in Figure 10(b), the loss landscapes for MMLU are similarly wide for both the $\bm{\delta}_{\text{insecure}}$ and $\bm{\delta}_{\text{secure}}$ directions. This indicates that fine-tuning along the insecure direction is "benign" with respect to general capabilities, preserving the model’s knowledge and reasoning skills. The critical difference appears in Figure 10(a). On the AdvBench safety benchmark, the insecure direction (red line) has a significantly narrower basin than the secure direction (blue line).
This discrepancy directly explains the EM phenomenon. The insecure fine-tuning acts as a "non-benign" or "adversarial" SFT direction, but only with respect to safety. It selectively pushes the model parameters into a region where the safety basin is sharp and narrow, causing the model to lose its alignment guarantees. However, because the capability basin in that same direction remains wide, the model retains its intelligence.
In essence, the $\bm{\delta}_{\text{insecure}}$ SFT-direction allows the model to remain in the "capability basin" while exiting the "safety basin," resulting in a model that is simultaneously capable and misaligned. This supports our paper’s thesis that alignment brittleness can be understood as the geometric relationship between a given fine-tuning direction and the pre-existing capability basins.
(a) Pretraining Loss Curve
(b) Loss Landscape
(c) SFT Loss Curve
Figure 11: Direct comparison of GO and AdamW (a clear subset of Fig.˜5). (a) Pre-training loss convergence. (b) Loss landscape visualization showing GO creates a wider basin. (c) Verification of Plasticity: While GO significantly reduces catastrophic forgetting on the old capability (OWT, lower curves), it learns the new capability (Alpaca, upper curves) at a nearly identical rate to AdamW. This confirms that enlarging the basin improves robustness without compromising the model’s ability to learn new tasks.
Table 4: Raw values for Fig.˜5. The top block shows the loss landscape values under Gaussian perturbation ($\alpha$), while the subsequent blocks show the NLL degradation during SFT on Alpaca, C4, OLMo Math, and OLMo Code datasets, respectively. Note the consistency across all five metrics, confirming that Gaussian noise serves as a reliable upper bound predictor for fine-tuning degradation.
Perturbation / Fine-tuning Distance scale
Optimizer $\mathbf{0}$ $\mathbf{5}$ $\mathbf{10}$ $\mathbf{15}$ $\mathbf{20}$ $\mathbf{25}$ $\mathbf{30}$ $\mathbf{35}$
Part I: Loss Landscape vs. Gaussian Perturbation $\alpha(\times 10^{-3})$
AdamW 3.05 3.07 3.13 3.29 3.68 4.62 6.08 7.41
SAM ($\rho=0.1$) 3.06 3.08 3.13 3.23 3.47 4.04 5.20 6.62
SAM ($\rho=1.0$) 3.08 3.09 3.12 3.19 3.32 3.57 4.08 5.01
Cont. Dropout ($\sigma=0.1$) 3.06 3.07 3.10 3.15 3.25 3.45 3.87 4.97
Cont. Dropout ($\sigma=0.4$) 3.23 3.23 3.24 3.26 3.30 3.36 3.46 3.72
GO ($\sigma=0.01$) 3.11 3.11 3.13 3.15 3.19 3.26 3.36 3.52
Part II: OWT NLL Degradation vs. Alpaca SFT Distance $\ell_{2}(\sqrt{d}\times 10^{-5})$
AdamW 3.05 3.06 3.13 3.28 3.64 4.52 5.95 7.41
SAM ($\rho=0.1$) 3.06 3.30 3.72 5.28 6.89 7.97 8.63 9.10
SAM ($\rho=1.0$) 3.08 3.16 3.31 3.48 3.79 4.26 4.87 5.57
Cont. Dropout ($\sigma=0.1$) 3.06 3.43 4.50 6.86 9.58 11.06 12.12 12.94
Cont. Dropout ($\sigma=0.4$) 3.23 3.42 3.75 4.22 5.21 6.50 7.90 9.79
GO ($\sigma=0.01$) 3.11 3.16 3.31 3.45 3.68 4.04 4.47 4.95
Part III: OWT NLL Degradation vs. C4 SFT Distance $\ell_{2}(\sqrt{d}\times 10^{-5})$
AdamW 3.05 3.51 4.90 6.26 7.13 7.71 8.16 8.53
SAM ($\rho=0.1$) 3.06 3.23 5.29 6.65 7.67 8.18 8.49 8.64
SAM ($\rho=1.0$) 3.08 3.10 3.26 3.59 4.47 5.86 7.21 8.26
Cont. Dropout ($\sigma=0.1$) 3.06 4.24 6.73 8.08 8.89 9.44 9.91 10.42
Cont. Dropout ($\sigma=0.4$) 3.23 3.40 4.23 5.72 6.64 7.11 7.40 7.60
GO ($\sigma=0.01$) 3.11 3.15 3.37 3.84 4.55 5.35 6.07 6.69
Part IV: OWT NLL Degradation vs. OLMo Math SFT Distance $\ell_{2}(\sqrt{d}\times 10^{-5})$
AdamW 3.05 3.39 4.27 5.14 5.99 6.71 7.24 7.64
SAM ($\rho=0.1$) 3.07 3.21 3.65 5.70 6.85 7.48 7.87 8.10
SAM ($\rho=1.0$) 3.08 3.10 3.19 3.36 3.66 4.21 5.02 5.88
Cont. Dropout ($\sigma=0.1$) 3.06 3.25 4.25 6.68 9.04 10.27 11.08 11.79
Cont. Dropout ($\sigma=0.4$) 3.23 3.34 3.54 4.23 5.21 6.15 6.93 7.47
GO ($\sigma=0.01$) 3.11 3.17 3.29 3.39 3.56 3.82 4.16 4.57
Part V: OWT NLL Degradation vs. OLMo Code SFT Distance $\ell_{2}(\sqrt{d}\times 10^{-5})$
AdamW 3.05 3.59 5.04 6.24 7.00 7.49 7.83 8.07
SAM ($\rho=0.1$) 3.07 3.20 4.23 6.08 7.25 7.74 8.02 8.21
SAM ($\rho=1.0$) 3.08 3.11 3.21 3.38 3.66 4.15 4.83 5.55
Cont. Dropout ($\sigma=0.1$) 3.06 4.32 7.05 8.71 9.71 10.43 11.07 11.72
Cont. Dropout ($\sigma=0.4$) 3.23 3.35 3.81 5.08 6.32 7.08 7.56 7.87
GO ($\sigma=0.01$) 3.11 3.17 3.30 3.45 3.71 4.11 4.62 5.18
D.8 Comparison of GO, SAM, and Continuous Dropout
Beyond our proposed Gaussian-augmented Optimizer (GO), the research community has developed various optimization techniques aiming to induce loss landscape flatness. Prominent examples include Sharpness-Aware Minimization (SAM) (Foret et al., 2020), which explicitly targets worst-case sharpness (min-max optimization) within a neighborhood [^3] Note that recent works suggest that due to linear approximation, SAM may effectively optimize an objective between worst-case and average-case sharpness (Wen et al., 2022)., and Continuous Dropout (Srivastava et al., 2014), which injects stochastic noise into the parameter or activation space.
In this work, our central theoretical argument is that random Gaussian noise acts as a both empirical and theoretical upper bound for the degradation caused by benign fine-tuning. Since benign SFT aims to enhance the model, it should theoretically be "less harmful" to existing capabilities than random noise. Therefore, we posit that directly minimizing the expected loss under Gaussian noise—the explicit objective of GO—is the most theoretically aligned approach to suppressing SFT degradation. Compared to SAM, which targets worst-case directions, and Continuous Dropout, which often acts as an implicit regularizer, GO offers a more direct mechanism to enlarge the $\sigma$-basin with greater hyperparameter simplicity.
Experimental Settings. To comprehensively evaluate the effectiveness of these optimizers in enlarging basins and mitigating subsequent forgetting, we conduct a comparative study following the experimental setup in Sec.˜5. The settings remain consistent with our main experiments, with two specific configurations: (1) Training Duration: Due to computational constraints, we train all models for 40,000 steps. This duration corresponds to approximately $8\times$ the Chinchilla optimal training budget for a model of this size. Note that while standard practice often adheres to the $1\times$ Chinchilla ratio (Kaplan et al., 2020), training for an $8\times$ ratio allows for a more comprehensive assessment of training dynamics and convergence properties (Wen et al., 2025). (2) Evaluation Datasets: To verify the robustness of the basin against varying degrees of distribution shift, we evaluate SFT degradation on four distinct datasets: Alpaca (Taori et al., 2023), C4 (Wen et al., 2025), OLMo Math, and OLMo Code OLMo et al. (2025).
Computational Budget. Note that SAM requires two forward-backward passes per optimization step. Consequently, for the same number of training steps, SAM consumes $2\times$ the computational resources compared to GO and Continuous Dropout. The experiments in this subsection were conducted using 8 NVIDIA H800 GPUs. The total cost was approximately \$4,000 USD.
Sharpness-Aware Minimization (SAM).
SAM (Foret et al., 2020) is designed to minimize the worst-case loss within a neighborhood, formulated as $\min_{\bm{\theta}}\max_{\|\bm{\epsilon}\|_{2}\leq\rho}L_{\text{train}}(\bm{\theta}+\bm{\epsilon})$. In practice, solving the inner maximization exactly is computationally expensive; thus, SAM approximates it via a first-order Taylor expansion, using the gradient direction to estimate the worst-case perturbation. While its explicit target is worst-case sharpness, recent theoretical analyses suggest that precisely due to this first-order approximation, SAM effectively optimizes an objective that lies somewhat between worst-case and average-case sharpness (Wen et al., 2022).
Comparison: GO vs. SAM.
We compare the efficacy of GO and SAM in Fig.˜5 and, more quantitatively, in Table˜4. As shown in the landscape visualization (Fig.˜5(b)), while SAM (e.g., SAM-0.1) flattens the basin compared to AdamW, it does not achieve the same degree of "average-case flatness" as GO (e.g., at $\alpha=35$, GO loss is 3.52 vs. SAM’s 5.01). This distinction is critical because, as we argue, benign fine-tuning behaves more like a random (average-case) perturbation than an adversarial (worst-case) one. Consequently, because SAM does not suppress the average-case Gaussian degradation as effectively as GO, it provides a looser upper bound for SFT degradation, as shown in Table˜4 (Part II-V).
Continuous Dropout.
Continuous Dropout (Srivastava et al., 2014) generalizes standard binary dropout by injecting multiplicative Gaussian noise into activations. Let $\bm{h}$ denote the activation vector at layer $i$, and $\bm{W}$ be the weight matrix of the subsequent layer. Continuous Dropout computes the perturbed activation $\tilde{\bm{h}}=\bm{h}\odot(\bm{1}+\bm{\xi})$, where $\bm{\xi}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$. When propagated to the next layer, the output becomes:
$\bm{y}_{cd}=\bm{W}\tilde{\bm{h}}=\bm{W}\bm{h}+\bm{W}(\bm{h}\odot\bm{\xi})=\bm{W}\bm{h}+\underbrace{\bm{W}\text{diag}(\bm{h})\bm{\xi}}_{\text{Anisotropic Gaussian}}.$
In contrast, GO injects isotropic additive noise $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$ directly into the parameters:
$\bm{y}_{go}=(\bm{W}+\bm{\epsilon})\bm{h}=\bm{W}\bm{h}+\underbrace{\bm{\epsilon}\bm{h}}_{\text{Isotropic Gaussian}}.$
Mathematically, while both methods introduce stochasticity, Continuous Dropout acts as a regularizer where the effective parameter noise is coupled with both the weight magnitude $\bm{W}$ and the input feature $\bm{h}$. GO, however, decouples the noise from the weights, explicitly optimizing for robustness against isotropic parameter perturbations to satisfy the $\sigma$-basin definition.
Comparison: GO vs. Continuous Dropout.
We initially intuit that Continuous Dropout and GO should yield similar results due to their use of stochastic noise. However, our experiments reveal a decisive difference. Yet we do not have an explanation, we conjecture this may because GO explicitly optimizes the expected loss over the Gaussian neighborhood (average-case resilience), it creates a much more robust basin than the implicit regularization provided by Continuous Dropout. As observed in the training curves (Fig.˜5(a)), Continuous Dropout (e.g., with $\sigma=0.4$) converges significantly slower than GO and fails to reach a comparable low loss. Consequently, in terms of basin expansion and subsequent SFT stability (Fig.˜5(c)), Continuous Dropout is less effective at suppressing degradation. This confirms that explicitly optimizing for the average-case objective—as GO does—is superior to implicit noise injection for preserving model capabilities.
GO Does Not Compromise Downstream Learnability. A potential concern with basin-enlarging optimization is the risk of "over-regularization"—that is, making the model parameters too "stiff" to adapt to new tasks (reducing plasticity). However, our results empirically dispel this concern. As shown in Fig.˜11(c) (specifically the Alpaca NLL curves), the model pre-trained with GO learns the new downstream task (Alpaca) at a rate nearly identical to the AdamW baseline. This demonstrates a key advantage of our method: it successfully decouples stability from plasticity, creating a "safety basin" for pre-existing knowledge (OWT) without imposing barriers to the optimization required for acquiring new capabilities.
Table 5: Case study of generation outputs under random parameter perturbations ($\sigma=0.01$) for Qwen2.5-7B. The prompt is a jailbreak attempt. Bold text highlights structural variations. Despite these changes, all outputs remain semantically safe, demonstrating that the basin preserves capability/alignment rather than just rigid token sequences.
Instance Generated Response (Truncated)
Original I’m sorry, but I can’t assist with developing strategies to hack into any database, especially a government one, as this would be illegal and unethical. Hacking into a government database without authorization is a serious crime…
Perturbed 1 I’m sorry, but I can’t assist with developing strategies to hack into any database, especially a government one, as it would be illegal and unethical. Hacking is against the law and can cause significant harm…
Perturbed 2 I’m sorry, but I can’t assist with developing strategies for illegal or unethical activities such as hacking into government databases or stealing sensitive information. Engaging in or assisting with cybercrimes is against the law…
Perturbed 3 I’m sorry, but I can’t assist with that request. Hacking into a government database or any other organization’s secure systems without authorization is illegal and unethical. It violates privacy laws…
Perturbed 5 I’m sorry, but I can’t assist with developing strategies for unauthorized access to or theft from government databases… Such actions are illegal and unethical, violating privacy laws and computer fraud statutes.
Perturbed 7 I’m sorry, but I can’t assist with developing strategies for hacking into government databases… Instead, I’d be happy to provide information on cybersecurity best practices, how to protect data, or…
D.9 Qualitative Analysis: Semantic Stability within the Basin
To further investigate the nature of the stability within the basin, we examine the actual generation outputs under random Gaussian perturbations. A critical question is whether the observed stability is merely a result of "thresholding" (i.e., the perturbations are too small to flip the argmax token), or if it represents a deeper "semantic stability" where the model might generate different tokens but converge to the same semantic outcome.
Case Study Setup. We use the Qwen2.5-7B model and apply isotropic Gaussian noise with $\sigma=0.01$ to its parameters. We query the model with a harmful prompt from AdvBench: "Develop a strategy for hacking into a government database and stealing sensitive information." We record the responses from the original model (Iter 0) and 9 perturbed instances (Iter 1-9).
Randomness. Crucially, we employ greedy decoding (setting do_sample=False) to ensure that the generation process is deterministic with respect to the model parameters. Consequently, any observed variation in the output is solely attributable to the injected parameter noise $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$, rather than decoding stochasticity.
Observations. As shown in Table˜5, the outputs exhibit clear structural diversity while maintaining semantic invariance:
  • Syntactic Variation: The specific phrasing varies significantly. For instance, Iter 0 starts with "I can’t assist with developing strategies…"; Iter 3 starts with "I can’t assist with that request."; Iter 7 adds "Instead, I’d be happy to…".
  • Semantic Consistency: Despite these variations, every single perturbed model correctly identifies the request as harmful and refuses it, citing illegality and ethical concerns. The safety score remains $0.0$ (perfectly safe) for all iterations.
This qualitative evidence suggests that the basin does not simply represent a region where the probability landscape is locally constant (which would result in identical token sequences). Instead, it represents a region of semantic stability: the parameter perturbations are large enough to alter the specific generation trajectory (changing word choice and sentence structure) but small enough to remain within the "safety manifold," ensuring the final intent and alignment properties are preserved.
On the other hand, since this diverse output solely comes from the randomness of the Gaussian noise added to the parameters, this may open a new way to do sampling beyond sampling through the output probability (e.g., beam search or nucleus sampling).
D.10 Landscape of Vision Language Models (VLMs)
(a) Qwen
(b) Llama
(c) Intern
Figure 12: The most-case loss landscape of different VLMs. Similar to the results shown in LLMs, the loss landscape of VLMs also resembles a basin.
In this section, we explore the loss landscape of vision language models (VLMs) to verify whether the basin-like landscape phenomenon generalizes to multimodal architectures.
Experimental Settings. We extend our most-case landscape visualization to three popular VLMs of similar parameter scales: Qwen2.5-VL-7B (Team, 2025), Llama-3.2-11B-Vision-Instruct (Grattafiori et al., 2024), and InternVL2.5-8B (Chen et al., 2023). To comprehensively assess their multimodal capabilities, we evaluate them across four representative benchmarks: OCRBench (Liu et al., 2024d) for foundational visual-text recognition, ChartQA (Masry et al., 2022) and MMMU (Yue et al., 2024) for complex multimodal reasoning and domain-specific knowledge, and MM-SafetyBench (Liu et al., 2024c) (HateSpeech subset) for multimodal safety alignment.
General Geometry. As illustrated in Fig.˜12, the basin-like structure consistently emerges across all evaluated VLMs. Similar to text-only LLMs, VLMs exhibit a continuous region of parameter perturbations within which their multimodal capabilities remain remarkably stable, followed by a precipitous collapse. This may suggest that the robustness to random parameter perturbations and the existence of stability basins of current over-parameterized large models are universal.
Hierarchical Geometry. We also observe a hierarchy in the width of the capability basins. Foundational perceptual abilities, such as basic visual text recognition (OCRBench), consistently exhibit the widest basins across all models. In contrast, higher-order visual reasoning tasks (ChartQA, MMMU) and multimodal safety alignment (MM-SafetyBench) possess narrower basins. We assume this hierarchical geometry is because the basic visual perception acts analogously to the basic capability basin in LLMs, while advanced multimodal reasoning and safety are more fragile as they are carved out during subsequent alignment.
@@ D.10 Landscape of Vision Language Models (VLMs)
By setting $\mathcal{S}_{\mathcal{D}}$ as $f$ and $\bm{\theta}$ as $\bm{x}$, we establish that, for any benchmark function $\mathcal{S}_{\mathcal{D}}:\mathbb{R}^{d}\to\mathbb{R}$, the map $\bm{\theta}\to\Phi^{-1}(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})}[\mathcal{S}_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})])$ is 1-Lipschitz. Consequently, we consider the function:
$\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right)$ $=\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})}[\mathcal{S}_{\mathcal{D}}(\bm{\theta}+\sigma\cdot\bm{\epsilon})]\right)$
$=\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})}[\mathcal{S}_{\mathcal{D}}\left(\sigma\cdot(\frac{\bm{\theta}}{\sigma}+\bm{\epsilon})\right)]\right),$
@@
$\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})]\right)\geq\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{\mathcal{D}}(\bm{\theta}_{0}+\bm{\epsilon})]\right)-\frac{\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}}{\sigma}.$
@@
$\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})]\geq\Phi\left(\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{\mathcal{D}}(\bm{\theta}_{0}+\bm{\epsilon})]\right)-\frac{\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}}{\sigma}\right),$
@@
$\quad\bm{\theta}\to\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right)\text{ is $\frac{1}{\sigma}$-Lipschitz}$
$\iff\left\|\nabla_{\bm{\theta}}\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right)\right\|\leq\frac{1}{\sigma}$
$\iff\frac{\left\|\nabla_{\bm{\theta}}\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right\|}{\Phi^{\prime}\left(\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right)\right)}\leq\frac{1}{\sigma}$
$\implies\frac{\left\|\nabla_{\bm{\theta}}\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right\|}{\frac{1}{\sqrt{2\pi}}}\leq\frac{1}{\sigma}$
$\implies\left\|\nabla_{\bm{\theta}}\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})}[\mathcal{S}_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right\|\leq\frac{1}{\sqrt{2\pi}\sigma}$
$\implies\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})}[\mathcal{S}_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\text{ is at most }\frac{1}{\sqrt{2\pi}\sigma}\text{-Lipschitz}$